y + 2y = 0 -Fina first three terms ies sln. two
Answer #1

n=0 リール nana Given equation is Y+2y=0 - fet y = {cath y = y - i From *, and e. We have E 2 (0-1) Gux n=1 n-2 2 n(n-1) Guxh